OpenStudy (1645323):

How to SImplify <-4,4>/sqrt32

OpenStudy (1645323):

@AndrewTheCookie

OpenStudy (jdoe0001):

\(\bf \cfrac{<-4,4>}{\sqrt{2}}?\)

OpenStudy (1645323):

yes

OpenStudy (jdoe0001):

hmm rather \(\bf \cfrac{<-4,4>}{\sqrt{32}}?\)

OpenStudy (1645323):

no sqrt 32 on bottom

OpenStudy (1645323):

yea that now how do i simpify that?

OpenStudy (jdoe0001):

looks like an scaler vector multiplication though => \(\bf \cfrac{<-4,4>}{\sqrt{32}}\implies <-4,4>\cfrac{1}{\sqrt{32}}\)

OpenStudy (1645323):

it is so do i multiply <-4,4> by sqrt32? so -4sqrt32 and factor 32 out?

OpenStudy (jdoe0001):

yeap.. scalar multiplication is just a<b,c> => <a*b, a*c>

OpenStudy (1645323):

ok

OpenStudy (jdoe0001):

keep in mind that \(\bf \sqrt{32}=\sqrt{2^5}\)